Write the equation in the slope-intercept form

m = 9 (3,1)

Found 2 solutions by Alan3354, Edwin McCravy:
Answer by Alan3354(69443)

Do it like this one:
--------------
Write the equation in the slope-intercept form
m = 4; (3,4)
---------------------
Use y = mx + b and the point to find b
4 = 4*3 + b
b = -8
--> y = 4x - 8

Answer by Edwin McCravy(20054)

Another way is to use the point-slope method.

y - y1 = m(x - x1)

Substituting m = 9, x1 = 3, y1 = 1

y - 1 = 9(x - 3)

y - 1 = 9x - 27

    y = 9x - 26

Edwin